# Using the antenna impedance to estimate soil electrical parameters for the MIST global 21-cm experiment

### v1

@@ Abstract
Context. Radio experiments trying to detect the global $21$-cm signal from the early Universe are very sensitive to the electrical properties of their environment. For ground-based experiments with the antenna above the soil it is critical to characterize the effect from the soil on the sky observations. This characterization requires estimating the soil’s electrical conductivity and relative permittivity in the same frequency range as the observations.
Aims. Here we present our initial effort to estimate the conductivity and relative permittivity of the soil using the impedance of an antenna mounted at a distance above the surface. The antenna used is part of the MIST global $21$-cm experiment.
Methods. We measured the antenna impedance at three sites in the Greater Concepción area, Chile. The measurements were done between $25$ and $125$ MHz, matching the range used by MIST for sky observations. The soil parameters were estimated by fitting the impedance measurements with electromagnetic simulations of the antenna and soil. In this initial effort the soil was modeled as homogeneous.
Results. The conductivity at the three sites was found to be between $0.007$ and $0.049$ Sm-1, and the relative permittivity between $1.6$ and $12.7$. The percent precision of the estimates at $68\%$ probability is, with one exception, better (lower) than $33\%$. The best-fit simulations have a better than $10\%$ agreement with the measurements relative to the peak values of the resistance and reactance across our frequency range. For MIST, these results represent a successful proof of concept of the use of the antenna impedance for soil characterization, and are expected to significantly improve in future implementations.
Key Words.:
Astronomical instrumentation, methods and techniques – Cosmology: observations, dark ages, first stars
@@ 1 Introduction
Ground-based experiments trying to detect the sky-averaged, or global, $21$-cm signal from the early Universe are sensitive to the electrical properties of the soil at the observation site (Bradley et al. 2019; Spinelli et al. 2022; Monsalve et al. 2024a; Pattison et al. 2024). The global $21$-cm signal, produced during the Universe’s dark ages, cosmic dawn, and epoch of reionization, should be detectable at frequencies 5similar-toabsent5$\sim$ 5∼ 5200200200200 MHz (Pritchard & Loeb 2012). At these frequencies, the system temperature of radio receivers is dominated by the contribution from astrophysical foregrounds, which is at least four orders of magnitude larger than the $21$-cm signal (Helmboldt & Kassim 2009; Guzmán et al. 2011; Panchenko et al. 2013; de Gasperin et al. 2020; Sasikumar Raja et al. 2022). Due to this high system temperature, detecting the global $21$-cm signal requires the calibration of the experiment’s instrumental response, including effects from the soil, with an accuracy of at least one part in $10^{4}$.
@@
Figure 1: MIST blade dipole antenna used to demonstrate the soil characterization technique. The antenna consists of two aluminum panels of size $1.2$ m $\times$ $60$ cm $\times$ $3$ mm, with a horizontal separation of $2$ cm. The panels are supported at a height of $50$ cm by a structure made out of fiberglass and plastic. The impedance of the antenna is measured in the gap between the two panels using an adapter. The adapter consists, at one end, of two copper wires bolted to each of the antenna panels and, at the other end, a female SMA connector. A $20$-m coaxial cable with SMA connectors is used to measure the antenna impedance with a VNA.
@@
Several ground-based global $21$-cm experiments, such as EDGES (Bowman et al. 2018), PRIZM Philip et al. (2019), LEDA (Spinelli et al. 2021), REACH (de Lera Acedo et al. 2022), SARAS (Singh et al. 2022), MIST (Monsalve et al. 2024a), and RHINO (Bull et al. 2024) have been designed around a single, wide-beam antenna. When this type of antenna operates close to the soil relative to the observation wavelength, the electromagnetic parameters, including the impedance and beam, are directly impacted by the soil’s electrical properties (Smith-Rose & Appleton 1933; Smith-Rose 1935). To reduce the effect of the soil on the antenna performance, EDGES, PRIZM, LEDA, and REACH, have opted for operating their antennas above a metal ground plane. In the case of EDGES, PRIZM, and LEDA, the ground plane is resting on the soil, while for REACH it is lifted above the soil (de Lera Acedo et al. 2022). Although finite metal ground planes do reduce the overall sensitivity of the antenna to the soil, they could also introduce artifacts into the sky measurements that are hard to estimate and remove (Bradley et al. 2019; Mahesh et al. 2021; Spinelli et al. 2022; Rogers et al. 2022). To avoid negative effects from metal ground planes, MIST has taken the approach of conducting sky observations without a ground plane (Monsalve et al. 2024a). This choice has the trade-off of folding the full effect from the soil into MIST’s electromagnetic performance. In this context, for MIST it becomes paramount to accurately determine the soil electrical parameters and integrate them into the analysis pipelines used for calibrating the sky observations (Monsalve et al. 2024b).
@@
In this paper, we demonstrate a technique to characterize the soil at the observation sites using the impedance of the MIST antenna. The parameters determined by this technique are the soil electrical conductivity and relative permittivity. These parameters are estimated by fitting measurements of the impedance with models from electromagnetic simulations of the antenna and soil. Our impedance measurements and simulations span the frequency range 25252525125125125125 MHz, matching the range used by MIST for observations of the sky spectrum (Monsalve et al. 2024a).
@@
Measurements of the antenna impedance are an essential part of MIST’s calibration program. They are conducted to calibrate the impedance mismatch between the antenna and the receiver input. However, here we explore the electrical characterization of the soil, also fundamental for MIST, as an additional use of the antenna impedance. Characterizing the soil using the antenna impedance could eliminate the need for additional measurements and equipment considered for this task.
@@
As indicated in IEEE (2011), using the impedance of an antenna located above the surface is a valid non-invasive approach for the electrical characterization of the soil at radio frequencies. This technique has been used for decades to characterize the electrical properties of materials in an antenna’s vicinity. Early results were obtained by comparing the measured impedance with analytical models (e.g., Bhattacharyya 1963; Wait 1969; Smith & King 1974; Wong et al. 1977; Nicol & Ridd 1988; He & Shen 1992). In several recent efforts the measurements have been compared with models from electromagnetic simulations (e.g., Wakita et al. 2000; Lenler-Eriksen & Meincke 2004; Berthelier et al. 2005; Le Gall et al. 2006). Our implementation of this technique represents an improvement on the latter works thanks to the availability of more powerful computers.
@@ 2.1 Blade dipole antenna
Panel (a) of Fig. 1 shows the blade dipole antenna used in this work. This antenna is a replica of the antennas used by MIST for the measurements described in Monsalve et al. (2024a). The antenna consists of two solid aluminum panels of $1.2$ m in length and $60$ cm in width, mounted horizontally $50$ cm above the soil. The panel thickness is $3$ mm and the horizontal separation between the panels is $2$ cm. The antenna is supported by a structure made out of fiberglass tubes and angles, as well as plastic connectors, screws, and nuts.
The antenna is fed through an adapter located in the gap between the panels. This adapter is shown in panel (b) of Fig. 1. The adapter consists, at one end, of two copper wires bolted to each of the antenna panels and, at the other end, a female subminiature version A (SMA) connector. The antenna impedance is measured using a coaxial cable connected to the adapter. The adapter does not provide direct current (DC) isolation between the panels and the cable. Specifically, one panel has DC continuity with the center conductor of the cable and the other panel has DC continuity with the outer conductor. As it leaves the antenna, the cable runs in the middle of the gap between the panels and perpendicular to the dipole excitation axis.
The full MIST instrument also includes a receiver box installed under the antenna, which contains the electronics and batteries, and a balun connected between the antenna and the receiver box (Monsalve et al. 2024a). In this initial study, however, we test the soil characterization technique using the impedance of the antenna by itself, without the receiver box and balun.
@@ 2.2 Test sites
Measurements were conducted at three sites in the Greater Concepción area, Chile, on November 14, 2023. The sites are shown in Fig. 2 and their coordinates are listed in Table 1. Sites A, B, and C, are located on a beach, a football field, and a park, respectively. These sites were selected because they have similar characteristics to the sites used by MIST for sky observations (Monsalve et al. 2024a). At the three sites the surface is level to within a few degrees and relatively flat, with vertical variations within $\sim 5$ cm ($25$ cm) inside a $\sim 2$-m ($20$-m) radius from the antenna. Within a $\sim 20$-m radius, there are no tall plants, trees, or buildings that could cause signal reflections. At Site A, the antenna is also more than $20$ m away from the sea water. Table 1 gives a general description of the soil composition at the surface.
@@ 2.3 Measurement setup
The measurements were done with a NanoVNA-H [^1] https://nanovna.com/ vector network analyzer (VNA), in the frequency range 25252525125125125125 MHz, with a $1$-MHz resolution. The data provided by the VNA correspond to the complex reflection coefficient, $\Gamma$, which we convert to impedance, $Z$, using the following equation (Pozar 2005):
@@
We connected the VNA to the antenna using a $20$-m RG174 cable with male SMA connectors. The VNA was operated manually and the operator was located by the VNA, $\sim 20$ m away from the antenna. The measurements were done with the operator sitting on the ground to minimize signal blockages and reflections. Before each antenna measurement, the setup was calibrated using the generic open, short, and load (OSL) SMA standards included with the NanoVNA-H. The calibration was done at the antenna end of the cable and, therefore, removed the unwanted S-parameters of the VNA and cable.
@@ 2.4 Measurements
The measurements are shown in Fig. 3. Panels (a) and (b) show, respectively, the magnitude and phase of the reflection coefficient. Panels (c) and (d) show, respectively, the resistance ($R$, real part of the impedance) and reactance ($X$, imaginary part of the impedance). The differences observed between the three sites are primarily due to differences in the electrical properties of the soils. The common features across sites include: (1) a dip in $|\Gamma|$ at $\sim 50$ MHz; (2) a fast transition in $\angle\Gamma$ at $\sim 50$ MHz; (3) a peak in $R$ at 65similar-toabsent65$\sim$ 65∼ 6575757575 MHz; and (4) a fast transition in $X$ at 60similar-toabsent60$\sim$ 60∼ 6080808080 MHz. Similar features are observed in the measurements done by MIST in California, Nevada, and the Canadian High Arctic (Monsalve et al. 2024a).
The measurement noise in the resistance and reactance has a sample standard deviation $<0.2$ $\Omega$, which was estimated from residuals resulting from polynomial fits to the data. For this initial study we did not estimate the systematic uncertainty in the measurements, which is primarily associated with the accuracy and stability of the OSL calibration. We expect this systematic uncertainty to dominate the measurement uncertainty budget. The determination of this uncertainty will be one of our priorities in future implementations of this technique.
Figure 4: Geometry of the Feko simulations. The simulations included the soil, $20$-m cable, antenna panels, and wires and SMA connector in the gap between the panels. The soil was simulated as homogeneous, and infinite in the horizontal directions and in depth. The meshing seen on the geometry (except on the soil) corresponds to a discretization of the surfaces for the computation of currents and fields with the method of moments.
@@ 3 Electromagnetic simulations
The electromagnetic simulations of the antenna were computed with the Feko software222www.altair.com/feko and its implementation of the method of moments. Figure 4 shows the geometry of the simulations. The geometry was designed to closely match the real antenna in order to maximize the simulation accuracy. Details of the simulation geometry are provided in Appendix A.

### v2

# Using the Antenna Impedance to Estimate Soil Electrical Parameters for the MIST Global 21 cm Experiment

@@ Abstract
Radio experiments trying to detect the global $21$ cm signal from the early Universe are very sensitive to the electrical properties of their environment. For ground-based experiments with the antenna above the soil it is critical to characterize the effect from the soil on the sky observations. This characterization requires estimating the soil’s electrical conductivity and relative permittivity in the same frequency range as the observations. Here we present our initial effort to estimate the conductivity and relative permittivity of the soil using the impedance of an antenna mounted at a distance above the surface. In this technique, the antenna used for soil characterization is the same as the antenna used for sky observations. To demonstrate the technique we use the antenna of the MIST global $21$ cm experiment. We measured the antenna impedance at three sites in the Greater Concepción area, Chile. The measurements were done between $25$ and $125$ MHz, matching the range used by MIST for sky observations. The soil parameters were estimated by fitting the impedance measurements with electromagnetic simulations of the antenna and soil. In this initial effort the soil was modeled as homogeneous. The conductivity at the three sites was found to be between $0.007$ and $0.049$ Sm-1, and the relative permittivity between $1.6$ and $12.7$. The percent precision of the estimates at $68\%$ probability is, with one exception, better (lower) than $33\%$. The best-fit simulations have a better than $10\%$ agreement with the measurements relative to the peak values of the resistance and reactance across our frequency range. For MIST, these results represent a successful proof of concept of the use of the antenna impedance for soil characterization, and are expected to significantly improve in future implementations.
Intergalactic medium (813); Calibration (2179); Radio receivers (1355); Early universe (435); Astronomical instrumentation (799)
software: SciPy (Virtanen et al., 2020)
[…]
2025 August 29
@@ 1 Introduction
Ground-based experiments trying to detect the sky-averaged, or global, $21$ cm signal from the early Universe are sensitive to the electrical properties of the soil at the observation site (Bradley et al., 2019; Mahesh et al., 2021; Spinelli et al., 2022; Monsalve et al., 2024a, b; Pattison et al., 2025). The global $21$ cm signal, produced during the Universe’s dark ages, cosmic dawn, and epoch of reionization, should be detectable at frequencies $\sim$55200200 MHz (Pritchard & Loeb, 2012). At these frequencies, the system temperature of radio receivers is dominated by the contribution from astrophysical foregrounds, which is at least four orders of magnitude larger than the $21$ cm signal (Helmboldt & Kassim, 2009; Guzmán et al., 2011; Panchenko et al., 2013; de Gasperin et al., 2020; Sasikumar Raja et al., 2022). Due to this high system temperature, detecting the global $21$ cm signal requires the calibration of the experiment’s instrumental response, including effects from the soil, with an accuracy of at least one part in $10^{4}$.
@@
Figure 1: MIST blade dipole antenna used to demonstrate the soil characterization technique. The antenna consists of two aluminum panels of size $1.2$ m $\times$ $60$ cm $\times$ 33 mm, with a horizontal separation of 22 cm. The panels are supported at a height of $50$ cm by a structure made out of fiberglass and plastic. The impedance of the antenna is measured in the gap between the two panels using an adapter. The adapter consists, at one end, of two copper wires bolted to each of the antenna panels and, at the other end, a female SMA connector. A 2020 m coaxial cable with SMA connectors is used to measure the antenna impedance with a VNA.
@@
Several ground-based global $21$ cm experiments, such as EDGES (Bowman et al., 2018), PRIZM (Philip et al., 2019), LEDA (Spinelli et al., 2021), REACH (de Lera Acedo et al., 2022), SARAS (Singh et al., 2022), MIST (Monsalve et al., 2024a), and RHINO (Bull et al., 2024), have been designed around a single, wide-beam antenna. When this type of antenna operates close to the soil relative to the observation wavelength, the electromagnetic parameters, including the impedance and beam, are directly impacted by the soil’s electrical properties (Smith-Rose & Appleton, 1933; Smith-Rose, 1935). To reduce the effect of the soil on the antenna performance, EDGES, PRIZM, LEDA, and REACH, have opted for operating their antennas above a metal ground plane. In the case of EDGES, PRIZM, and LEDA, the ground plane is resting on the soil, while for REACH it is lifted above the soil (de Lera Acedo et al., 2022).
Although finite metal ground planes significantly reduce the sensitivity of the antenna to the soil, they can also introduce frequency dependence, or chromaticity, into the antenna beam, which is typically unwanted. This chromaticity is produced by signal reflections from the edges of the ground plane, which represent an electrical discontinuity between the metal and the soil (Rogers et al., 2022). For the ground plane sizes used by current instruments, of order tens of meters on a side, this beam chromaticity generates spectral structure in the measured antenna temperature on scales comparable to global $21$ cm signal models (Pritchard & Loeb, 2012; Mahesh et al., 2021; Spinelli et al., 2022). If uncorrected or imperfectly accounted for, this structure could be confused for the cosmological signal or complicate its extraction. Furthermore, problems can occur due to imperfections in the ground plane such as unintended slots in the metal, which would generate resonances (Bradley et al., 2019), and bumps, which would produce additional reflections (Rogers et al., 2022). These imperfections can be difficult to diagnose and their effects difficult to estimate and remove during data analysis. Finally, since finite metal ground planes do not eliminate the sensitivity to the soil, it is necessary to characterize the soil and account for its impact even if a ground plane is used (Monsalve et al., 2017; Bowman et al., 2018; Mahesh et al., 2021; Monsalve et al., 2021; de Lera Acedo et al., 2022; Spinelli et al., 2022; Pattison et al., 2025).
To avoid the negative effects mentioned above, MIST has taken the approach of observing the sky with the antenna above the soil but without using a metal ground plane (Monsalve et al., 2024a). This approach has the trade-off of folding the full effect from the soil into MIST’s electromagnetic performance. Although the magnitude of the soil effect is larger in an absolute sense, it is possible that the spectral scales affected in this approach represent a more favorable scenario than when using a metal ground plane. Whether or not this approach yields better results will depend on the specific characteristics of the soil and how we account for its effects during analysis. In our implementation, the effect from the soil on the antenna impedance is directly determined from the impedance measurements conducted in the field (Monsalve et al., 2024a). On the other hand, the effect on the beam directivity and antenna losses is estimated using electromagnetic simulations of the instrument that include the soil. For these simulations, it is critical to have accurate knowledge of the soil’s electrical characteristics.
In Singh et al. (2022) SARAS measured the sky from a lake, thus avoiding the use of a large metal ground plane similarly to MIST. Due to its homogeneity, water is an ideal medium to use as a “soil” in real observations as well as in the context of simulations. However, operating over a sufficiently large body of water poses serious logistical challenges. These challenges are increased if the observation site is in a remote radio-quiet location. In contrast, operating above soil typically provides the safe and stable conditions required for the careful study of systematic effects and for achieving a high signal-to-noise ratio in the sky observations.
To maximize the accuracy of the electromagnetic simulations, MIST’s observation sites are chosen optimizing for surface flatness and homogeneous composition. Although inhomogeneities are unavoidable in natural soils, their effect might become subdominant when averaged over hundreds of meters around the antenna. In this case, simulating the soil as homogeneous or stratified could represent a sufficiently accurate approximation. This possibility is one of MIST’s research hypotheses. Leveraging MIST’s small size and portability, our experimental program includes operating from different sites to improve the understanding and mitigation of soil effects (Monsalve et al., 2024a).
@@
In this paper, we demonstrate a technique for the electrical characterization of the soil at sites used for observations of the global $21$ cm signal. In this technique, the electrical properties of the soil are inferred from the impedance of the antenna used for sky observations. The parameters determined by this technique are the soil’s bulk electrical conductivity and relative permittivity. These parameters are estimated by fitting impedance measurements with impedance models obtained from electromagnetic simulations of the antenna and soil.
@@
We demonstrate the technique using the antenna from the MIST experiment. Our impedance measurements and simulations span the frequency range 2525125125 MHz, matching the range used by MIST for observations of the sky spectrum (Monsalve et al., 2024a). Measurements of the antenna impedance are an essential part of MIST’s calibration program. They are conducted to calibrate the impedance mismatch between the antenna and the receiver input. However, here we explore the electrical characterization of the soil, also fundamental for MIST, as an additional use of the antenna impedance. Characterizing the soil using the antenna impedance could eliminate the need for additional measurements and equipment considered for this task.
@@
As indicated in IEEE (2011), using the impedance of an antenna located above the surface is a valid non-invasive approach for the electrical characterization of the soil at radio frequencies. This technique has been used for decades to characterize the electrical properties of materials in an antenna’s vicinity. Early results were obtained by comparing the measured impedance with analytical models (e.g., Bhattacharyya, 1963; Wait, 1969; Smith & King, 1974; Wong et al., 1977; Nicol & Ridd, 1988; He & Shen, 1992). In several recent efforts the measurements have been compared with models from electromagnetic simulations (e.g., Wakita et al., 2000; Lenler-Eriksen & Meincke, 2004; Berthelier et al., 2005; Le Gall et al., 2006). Our implementation of this technique represents an improvement on the latter works thanks to the availability of more powerful computers.
@@ 2.1 Blade Dipole Antenna
Panel (a) of Figure 1 shows the blade dipole antenna used in this work. This antenna is a replica of the antennas used by MIST for the sky measurements described in Monsalve et al. (2024a). The antenna consists of two solid aluminum panels of $1.2$ m in length and $60$ cm in width, mounted horizontally $50$ cm above the soil. The panel thickness is $3$ mm and the horizontal separation between the panels is $2$ cm. The antenna is supported by a structure made out of fiberglass tubes and angles, as well as plastic connectors, screws, and nuts.
The antenna is fed through an adapter located in the gap between the panels. This adapter is shown in panel (b) of Figure 1. The adapter consists, at one end, of two copper wires bolted to each of the antenna panels and, at the other end, a female subminiature version A (SMA) connector. The antenna impedance is measured using a coaxial cable connected to the adapter. The adapter does not provide direct current (DC) isolation between the panels and the cable. Specifically, one panel has DC continuity with the center conductor of the cable and the other panel has DC continuity with the outer conductor. As it leaves the antenna, the cable runs in the middle of the gap between the panels and perpendicular to the dipole excitation axis.
The full MIST instrument also includes a receiver box installed under the antenna, which contains the electronics and batteries, and a balun connected between the antenna and the receiver box (Monsalve et al., 2024a). In this initial study, however, we test the soil characterization technique using the impedance of the antenna by itself, without the receiver box and balun.
@@ 2.2 Test Sites
Measurements were conducted at three sites in the Greater Concepción area, Chile, on 2023 November 14. The sites are shown in Figure 2 and their coordinates are listed in Table 1. Sites A, B, and C, are located on a beach, a football field, and a park, respectively. These sites were selected because they have similar characteristics to the sites used by MIST for sky observations (Monsalve et al., 2024a). At the three sites the surface is level to within a few degrees and relatively flat, with vertical variations within $\sim$55 cm ($25$ cm) inside a $\sim$22 m ($20$ m) radius from the antenna. Within a $\sim$2020 m radius, there are no tall plants, trees, or buildings that could cause signal reflections. At Site A, the antenna is also more than $20$ m away from the sea water. Table 1 gives a general description of the soil composition at the surface.
@@ 2.3 Measurement Setup
The measurements were done with a NanoVNA-H [^1] https://nanovna.com/ vector network analyzer (VNA), in the frequency range 2525125125 MHz, with a $1$ MHz resolution. The data provided by the VNA correspond to the complex reflection coefficient, $\Gamma$, which we convert to impedance, $Z$, using the following equation (Pozar, 2005):
@@
We connected the VNA to the antenna using a $20$ m RG174 cable with male SMA connectors. The VNA was operated manually and the operator was located by the VNA, $\sim$2020 m away from the antenna. The measurements were done with the operator sitting on the ground to minimize signal blockages and reflections. Before each antenna measurement, the setup was calibrated using the generic open, short, and load (OSL) SMA standards included with the NanoVNA-H. The calibration was done at the antenna end of the cable and, therefore, removed the unwanted S-parameters of the VNA and cable.
@@ 2.4 Measurements
The measurements are shown in Figure 3. Panels (a) and (b) show, respectively, the magnitude and phase of the reflection coefficient. Panels (c) and (d) show, respectively, the resistance ($R$, real part of the impedance) and reactance ($X$, imaginary part of the impedance). The differences observed between the three sites are primarily due to differences in the electrical properties of the soils. The common features across sites include: (1) a dip in $|\Gamma|$ at $\sim$$50$ MHz; (2) a fast transition in $\angle\Gamma$ at $\sim$$50$ MHz; (3) a peak in $R$ at $\sim$65657575 MHz; and (4) a fast transition in $X$ at $\sim$60608080 MHz. Similar features are observed in the measurements done by MIST in California, Nevada, and the Canadian High Arctic (Monsalve et al., 2024a).
The measurement noise in the resistance and reactance has a sample standard deviation <0.2<0.2 $\Omega$, which was estimated from residuals resulting from polynomial fits to the data. For this initial study we did not estimate the systematic uncertainty in the measurements, which is primarily associated with the accuracy and stability of the OSL calibration. We expect this systematic uncertainty to dominate the measurement uncertainty budget. The determination of this uncertainty will be one of our priorities in future implementations of this technique.
Figure 4: Geometry of the Feko simulations. The geometry includes the soil, 2020 m cable, antenna panels, and wires and SMA connector in the gap between the panels. The soil was simulated as homogeneous, and infinite in the horizontal directions and in depth. The meshing seen in panel (b) on the geometry (except on the soil) corresponds to a discretization of the surfaces for the computation of currents and fields with the method of moments.
@@ 3 Electromagnetic Simulations
The electromagnetic simulations of the antenna were computed with the Feko software222www.altair.com/feko and its implementation of the method of moments. The geometry of the simulations is shown in Figure 4 and described in detail in Appendix A. The geometry includes the soil, the antenna panels, the wires and SMA connector that form the adapter in the gap between the panels, and the $20$ m cable starting in the gap and continuing toward the VNA. The geometry was designed using nominal dimensions of the antenna panels and $20$ m cable, and laboratory measurements of the adapter.
During a preliminary exploration, the geometry only consisted of the antenna panels, an idealized excitation port, and the soil. With this simple geometry, however, no soil parameter combinations could produce a reasonable agreement with the measurements (Appendix A.1). For this reason we incorporated the additional features (Appendix A.2), after which the agreement between the simulations and the data improved significantly.
The assembly of the antenna, adapter, and cable was replicated at the three test sites as closely as possible, and in this paper we use the same geometry for the simulations corresponding to the three sites. The simulations were conducted in the range 2525125125 MHz using double precision calculations and fine automatic mesh size.